Inbox (25) ale kscgi x ísl.exenou-igN slkr? 8 P 3 H-lvddbBx2ZxPuA yASwGHLApiAQHcF25VHHv5kUeqNRd Uk4 54 M 8 ym ?K ks.com er 2018 Ch 1 2 Ho The students of a certain college were asked to choose which of six movie genres was their favorite. The pie chart below shows the distribution of the students answers. If there are 13,500 students at this college, how many chose Drama or Historical? 2 dy
Answer #1

n=13500

Number of students who choose drama or historical=27% of n + 13% of n=27*13500/100+13*13500/100=5400

5400 choose drama or historical.